OpenStudy (sbentleyaz):

A particle moves along the x-axis so that at any time t, measured in seconds, its position is given by s(t) = sin(t) - 4cos(2t), measured in feet. What is the acceleration of the particle at time t = π seconds?

OpenStudy (steve816):

Hey! Welcome to OpenStudy!!!

OpenStudy (steve816):

You are looking for acceleration ya? Remember that acceleration is the second derivative of a position function. Can you find the second derivative of your equation?

OpenStudy (sbentleyaz):

I got -1 as the answer but i am not sure if its right

OpenStudy (sooobored):

the first derivative of the position function will yield the velocity function the 2nd derivative of the position function will yield the acceleration function or the first derivative of the velocity function s(t) = sin(t) - 4cos(2t) s'(t)= cos(t) +8sin(2t) s''(t) = -sin(t) + 16 cos(2t) if you want to find the acceleration when t= pi then you use the derived acceleration function s''(pi)=...

OpenStudy (sooobored):

comment- i did not get -1 as the acceleration

OpenStudy (sunnnystrong):

... ^ find the equation of the second derivative and plug in pi